OpenStudy (anonymous):

Bob paid $3. Less than Tim at a local pizza parlor. together they spent $7. 80. How much did Bob spend at the local pizza parlor?

OpenStudy (stefrheart):

x-3+x=7.80

OpenStudy (texaschic101):

B + T = 7 -- together they paid 7 B = T - 3 -- bob paid 3 less then tim now we sub T - 3 in for B in the first equation T - 3 + T = 7 -- combine like terms 2T - 3 = 7 -- add 3 to both sides 2T = 7 + 3 2T = 10 -- divide both sides by 2 T = 5 B + T = 7 B + 5 = 7 B = 7 - 5 B = 2 check.. B = T - 3 2 = 5 - 3 2 = 2 (correct) Bob spent $2 and Tim spent $5.....doesn't sound fair to me...lol